# A lower bound for classical Kloosterman sums and an application

### v3

@@ 3. An application
We denote $\lambda_{n}(f)$ to be the $n$-th normalised Hecke eigenvalue of $f$, i.e. $T_{n}(f)=\lambda_{n}(f)f$. The Fourier coefficient $a_{n}(f)$ and $\lambda_{n}(f)$ are related by the condition $a_{n}(f)=a_{1}(f)\lambda_{n}(f).$
@@
The above equation demonstrates how Petersson’s trace formula is indeed a weighted trace formula for Hecke operators. We use the following version of Petersson’s trace formula given in [ILS00, Proposition 2.1].
@@
$-\frac{3}{\nu}\leq J_{\nu}(\nu)-\frac{C_{0}}{\nu^{1/3}}\leq\frac{3}{\nu}.$
For all $\nu\geq\left(\frac{3}{\epsilon}\right)^{3/2},$ we have $\nu^{2/3}\geq\frac{3}{C_{0}-A(\epsilon)},$ i.e., $\frac{C_{0}}{\nu^{1/3}}-\frac{3}{\nu}\geq\frac{A(\epsilon)}{\nu^{1/3}}$ and this proves the lower bound for $J_{\nu}(\nu).$
@@
Since
@@
[…]
@@
Let $G(N)=\frac{f}{(\tau(f)^{2}f)^{\phi(f)/2\tau(f)}},$ for $N=df.$ Then, Corollary 4 gives
$\left|S(m,n;N)\right|\geq 2^{\omega(N)}\frac{G(N)}{N^{1/2}},$
Combining the previous equation with equation 3.11, we get
(3.12) $\begin{split}&\left\lvert\frac{S(m,n;N)}{N}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{N}\right)\right\rvert\geq 2^{\omega(N)}\frac{G(N)}{N^{3/2}}D_{0}^{k-1}J_{k-1}(k-1).\end{split}$
@@
i.e., $\left(D_{0}e^{-H_{0}}\right)^{k-1}\geq\frac{7}{H(N)},$ and this condition implies that
@@
$\begin{split}\frac{|B|}{2\pi}\leq 3.005e^{H_{0}(k-1)}J_{k-1}(k-1)\leq H(N)D_{0}^{k-1}J_{k-1}(k-1)\leq\left\lvert\frac{S(m,n;N)}{N}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{N}\right)\right\rvert.\end{split}$
@@
$\begin{split}&\left|\Delta_{k,N}(m,n)-\delta(m,n)\right|\\ =&\left|2\pi i^{k}\frac{S(m,n;N)}{N}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{N}\right)+B\right|\\ \geq&2\pi J_{k-1}(k-1)\left|H(N)D_{0}^{k-1}-3.005e^{H_{0}(k-1)}\right|\\ \geq&7.99\pi J_{k-1}(k-1)e^{H_{0}(k-1)},\end{split}$
@@
Applying Corollary 1 with its hypotheses, $\epsilon=0.4470,$ $D_{0}=0.999,$ $k\geq\max(22,k_{0}(N))$ and $\frac{4\pi\sqrt{mn}}{N}\in\left[0.999(k-1),(k-1)\right],$ we get
@@
The following corollaries are immediate. Although Corollary 5, Corollary 6 are mentioned for a specific value of $D_{0}=0.999,$ these results hold for any $D_{0}\in(e^{H_{0}},1),$ where $H_{0}$ is mentioned in Theorem 2.
@@
Let $k,\,N$ be natural numbers with $k\geq\max(22,k_{0}(N)),$ $k$ even and $N$ odd. Let $n\geq 2$ be a natural number such that $n\in\left[\left(\frac{0.999(k-1)N}{4\pi}\right)^{2},\left(\frac{(k-1)N}{4\pi}\right)^{2}\right]$ with $(n,N)=1$, and $n$ is a quadratic residue modulo each prime dividing the powerful part of $N$. Then we have
@@
Keeping equation (3.2) in mind, we note that Corollary 5 gives us a lower bound for a weighted trace of Hecke operator $T_{n}.$
Corollary 6.
Let $k,\,N$ be natural numbers with $k\geq\max(22,k_{0}(N)),$ $k$ even and $N$ odd. Let $n\geq 2$ be a natural number such that $n\in\left[\left(\frac{0.999(k-1)N}{4\pi}\right)^{2},\left(\frac{(k-1)N}{4\pi}\right)^{2}\right]$ with $(n,N)=1$, and $n$ is a quadratic residue modulo each prime dividing the powerful part of $N$. Then for $t>1,$ we have
$\sum_{f\in\mathcal{F}_{k}(N)}|\lambda^{t}_{n}(f)|\geq\left(\frac{(4\pi)^{k-1}}{\Gamma(k-1)}\right)^{t}\frac{1}{|\mathcal{F}_{k}(N)|^{t-1}}\left|\frac{0.002397\pi}{(k-1)^{1/3}}e^{(k-1)\left(1-\frac{4}{9}-\log(\frac{9}{5})\right)}\right|^{t}.$
Proof.
From equation (3.2), we have
$\Delta_{k,N}(1,n)=\sum_{f\in\mathcal{F}_{k}(N)}\left(\frac{\Gamma(k-1)}{(4\pi)^{k-1}}\right)|a_{1}(f)|^{2}\lambda_{n}(f).$
Using Hölder’s inequality for positive real number $t>1$ with $t,l:=\frac{t}{t-1}>1$ and Ramanujan-Deligne bound, we obtain
$|\Delta_{k,N}(1,n)|\leq\left(\frac{\Gamma(k-1)}{(4\pi)^{k-1}}\right)\left(\sum_{f\in\mathcal{F}_{k}(N)}|a_{1}(f)|^{l}\right)^{1/l}\left(\sum_{f\in\mathcal{F}_{k}(N)}|\lambda^{t}_{n}(f)|\right)^{1/t}$
$\leq\left(\frac{\Gamma(k-1)}{(4\pi)^{k-1}}\right)|\mathcal{F}_{k}(N)|^{1/l}\left(\sum_{f\in\mathcal{F}_{k}(N)}|\lambda^{t}_{n}(f)|\right)^{1/t}.$
Hence by Corollary 5, for any real number $t>1,$
$\sum_{f\in\mathcal{F}_{k}(N)}|\lambda^{t}_{n}(f)|\geq\left(\frac{(4\pi)^{k-1}}{\Gamma(k-1)}\right)^{t}\frac{1}{|\mathcal{F}_{k}(N)|^{t-1}}\left|\frac{0.002397\pi}{(k-1)^{1/3}}e^{(k-1)\left(1-\frac{4}{9}-\log(\frac{9}{5})\right)}\right|^{t}.$
Remark 6.
We note that the above corollary gives us a lower bound for the trace of the operator $T_{n}^{t}$ for an even positive integer $t>1$, whenever $n$ satisfies the hypothesis of Corollary 6.
@@
where $|\theta|\leq 1$ and $0\leq t\leq 1$ Note that for $k\geq 180,$
@@
$\begin{split}\frac{|B|}{2\pi}\leq 3.005e^{H_{0}(k-1)}J_{k-1}(k-1)\leq e^{H_{0}(k-1)}\frac{1.632}{(k-1)^{\frac{1}{3}}}\leq\frac{0.327H(N)}{(k-1)^{\frac{1}{3}}}\leq\left\lvert\frac{S(m,n;N)}{N}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{N}\right)\right\rvert.\end{split}$
@@
Note that the above theorem can be extended similarly to $\left[(k-1),(k-1)+\gamma(k-1)^{\frac{1}{3}}\right]$ with $-2^{\frac{1}{3}}\gamma>a$, where $a=-2.33811\dots$ is the largest zero of the airy function. This forces $\gamma<1.86.$ Furthermore, it is important to note that we can not apply [Kra14, Theorem 1] directly for the interval $\left[2,(k-1)\right)$ as [Kra14, Theorem 1] is valid for $t\geq 0.$
Proof of Corollary 3.
Using Hölder’s inequality for $t>1$, Ramanujan-Deligne bound and proceeding similarly as in the proof of Corollary 6, the proof is complete. ∎
@@
In this subsection, we consider some specific values of $N$ and compute the values of $H(N),k_{0}(N),k_{1}(N)$. Following Remark 4, we take $D_{0}=0.999$ so that $A_{0}=D_{0}e^{-H_{0}}\approx 1.0317$ as given in Theorem 2. We recall that $k_{0}(N)$ and $k_{1}(N)$ appear in Theorem 2 and Theorem 3 respectively. Recall that $N=df$ with $d$ is powerful and $f$ is squarefree, where $(d,f)=1$. We have $H(N)=2^{\omega(N)}\frac{G(N)}{N^{3/2}}$ with $G(N)=\frac{f}{(\tau(f)^{2}f)^{\phi(f)/2\tau(f)}}.$ If $N=p^{\alpha}$ for some prime $p$ and $\alpha>1$, then $H(N)=\frac{2}{p^{\frac{3\alpha}{2}}}.$ If $N$ is squarefree, then $H(N)=\frac{2^{\omega(N)}}{N^{1/2}(\tau(N)^{2}N)^{\phi(N)/2\tau(N)}}.$ If $N=p$ for a prime number, then $H(p)=\frac{2}{p^{1/2}(4p)^{(p-1)/4}}$.
We calculate the following table for $C_{0}=0.326$, $-H_{0}\geq 0.3223$ and $D_{0}=0.999.$ Then, $\log A_{0}\geq 0.031229.$
$N$ $\log H(N)(\geq)$ $k_{0}(N)=2+\left\lfloor\frac{\log 7-\log H(N)}{\log A_{0}}\right\rfloor(\leq)$ $k_{1}(N)=2+\left\lfloor\frac{\log(1.632/C_{0})-\log H(N)}{-H_{0}}\right\rfloor(\leq)$
$1$ 0 64 6
$3$1.0987 99 10
$5$3.1074 163 16
$9$2.6027 147 15
$49$5.1446 229 22
$99$12.5690 466 45
$999$71.6034 2357 229
$9999$1227.9235 39384 3816
$49999$152574.2718 4885724 473399
$99999$16325.0518 522817 50658
$999999$15030.9423 481377 46643
@@ Acknowledgements
The third named author acknowledges the support from Bhaskaracharya Pratisthan and IISER Pune. The second named author acknowledges the support of IISER Thiruvananthapuram. The second and third named authors would like to thank Kaneenika Sinha for her valuable input. We are thankful to the Harish-Chandra Research Institute, Prayagraj, for excellent hospitality while the parts of the work were carried out. We are grateful to the anonymous referee for pointing out an error in Theorem 1 and consequent results.

### v4

@@ 3. An application
We denote $\lambda_{n}(f)$ to be the $n$-th normalised Hecke eigenvalue of $f$, that is, $T_{n}(f)=\lambda_{n}(f)f$. The Fourier coefficient $a_{n}(f)$ and $\lambda_{n}(f)$ are related by the condition $a_{n}(f)=a_{1}(f)\lambda_{n}(f).$
@@
The above equation demonstrates how Petersson’s trace formula is indeed a weighted trace formula for Hecke operators.
We use the following version of Petersson’s trace formula given in [ILS00, Proposition 2.1].
@@
$\frac{3}{\nu}\leq J_{\nu}(\nu)-\frac{C_{0}}{\nu^{1/3}}\leq\frac{3}{\nu}.$
For all $\nu\geq\left(\frac{3}{\epsilon}\right)^{3/2},$ we have $\nu^{2/3}\geq\frac{3}{C_{0}-A(\epsilon)},$ that is, $\frac{C_{0}}{\nu^{1/3}}-\frac{3}{\nu}\geq\frac{A(\epsilon)}{\nu^{1/3}}$, and this proves the lower bound for $J_{\nu}(\nu).$
@@
Note that we have replaced $\log\frac{10}{9b}$ by $\log\frac{9b}{10}$. Since
@@
(3.8) $\begin{split}\sum_{b\geq 5}e^{-(k-1)\big(\frac{8}{9b}+\log\frac{9b}{10}\big)}\leq&\int_{4}^{\infty}e^{-(k-1)\big(\frac{8}{9x}+\log\frac{9x}{10}\big)}\,dx\\ =&\frac{25}{18}\int_{0}^{\frac{2}{9}}e^{-(k-1)y}\left(\frac{5y}{4}\right)^{k-3}\,dy\\ \leq&\frac{25}{18}\int_{0}^{\frac{2}{9}}\left(\frac{5}{18}\right)^{k-3}e^{-(k-1)y}\,dy\\ =&\frac{25}{18}\left(\frac{5}{18}\right)^{k-3}\left(\frac{1-e^{\frac{-2(k-1)}{9}}}{(k-1)}\right)\leq\frac{25}{18}\left(\frac{5}{18}\right)^{k-3}\frac{1}{(k-1)}.\end{split}$
@@
Let $G(N)=\frac{u}{(\tau(u)^{2}u)^{\phi(u)/2\tau(u)}},$ for $N=du.$ Then, Corollary  2 gives
$\left|S(m,n;N)\right|\geq 2^{\omega(N)}\frac{G(N)}{N^{1/2}}.$
Combining the previous equation with Equation (3.11), we get
(3.12) $\begin{split}&\left\lvert\frac{S(m,n;N)}{N}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{N}\right)\right\rvert\geq 2^{\omega(N)}\frac{G(N)}{N^{3/2}}D_{0}^{k-1}J_{k-1}(k-1)=H(N)D_{0}^{k-1}J_{k-1}(k-1).\end{split}$
@@
that is, $\left(D_{0}e^{-H_{0}}\right)^{k-1}\geq\frac{7}{H(N)},$ and this condition implies that
@@
[…]
@@
$\begin{split}\left|\Delta_{k,N}(m,n)-\delta(m,n)\right|=&\left|2\pi i^{k}\frac{S(m,n;N)}{N}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{N}\right)+B\right|\\ \geq&2\pi\left(\left|\frac{S(m,n;N)}{N}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{N}\right)\right|-\frac{|B|}{2\pi}\right)\\ \geq&7.99\pi J_{k-1}(k-1)e^{H_{0}(k-1)},\end{split}$
@@
Applying Corollary  1 with its hypotheses, $\epsilon=0.4470,$ $D_{0}=0.999,$ $k\geq\max(22,k_{0}(N))$ and $\frac{4\pi\sqrt{mn}}{N}\in\left[0.999(k-1),(k-1)\right],$ we get
@@
The following corollary is immediate. Although Corollary  3 is mentioned for a specific value of $D_{0}=0.999,$ this result holds for any $D_{0}\in(e^{H_{0}},1),$ where $H_{0}$ is mentioned in Theorem 2.
@@
Let k,Nk,\,N be natural numbers with $\geq\max(22,k_{0}(N)), kk$ even and $N$ odd. Let $n\geq 2$ be a natural number such that $n\in\left[\left(\frac{0.999(k-1)N}{4\pi}\right)^{2},\left(\frac{(k-1)N}{4\pi}\right)^{2}\right]$ with $(n,N)=1$, and $n$ is a quadratic residue modulo each prime dividing the powerful part of $N$. Then we have
@@
Keeping equation (3.2) in mind, we note that Corollary  3 gives us a lower bound for a weighted trace of Hecke operator $T_{n}.$
@@
where $|\theta|\leq 1$ and $0\leq t\leq 1.$ Note that for $k\geq 180,$
@@
[…]
@@
Note that the above theorem can be extended similarly to $\left[(k-1),(k-1)+\gamma(k-1)^{\frac{1}{3}}\right]$ with $2^{\frac{1}{3}}\gamma>a$, where $a=-2.33811\dots$ is the largest zero of the airy function. This forces $\gamma<1.86.$ Furthermore, it is important to note that we can not apply [Kra14, Theorem 1] directly for the interval $\left[2,(k-1)\right)$ as [Kra14, Theorem 1] is valid for $t\geq 0.$
@@
In this subsection, we consider some specific values of $N$ and compute the values of $H(N),k_{0}(N),k_{1}(N)$. Following Remark 4, we take $D_{0}=0.999$ so that $A_{0}=D_{0}e^{-H_{0}}\approx 1.0317$ as given in Theorem 2. We recall that $k_{0}(N)$ and $k_{1}(N)$ appear in Theorem 2 and Theorem 3 respectively. Recall that $N=du$ with $d$ is powerful and $u$ is squarefree, where $(d,u)=1$. We have $H(N)=2^{\omega(N)}\frac{G(N)}{N^{3/2}}$ with $G(N)=\frac{u}{(\tau(u)^{2}u)^{\phi(u)/2\tau(u)}}.$ If $N=p^{\alpha}$ for some prime $p$ and $\alpha>1$, then $H(N)=\frac{2}{p^{\frac{3\alpha}{2}}}.$ If $N$ is squarefree, then $H(N)=\frac{2^{\omega(N)}}{N^{1/2}(\tau(N)^{2}N)^{\phi(N)/2\tau(N)}}.$ If $N=p$ for a prime number, then $H(p)=\frac{2}{p^{1/2}(4p)^{(p-1)/4}}$.
We calculate the following table for $C_{0}=0.326$, $H_{0}\geq 0.3223$ and $D_{0}=0.999.$ Then, $\log A_{0}\geq 0.031229.$
NN $\log H(N)(\geq)$ $k_{0}(N)=2+\left\lfloor\frac{\log 7-\log H(N)}{\log A_{0}}\right\rfloor(\leq)$ $k_{1}(N)=2+\left\lfloor\frac{\log(1.632/C_{0})-\log H(N)}{-H_{0}}\right\rfloor(\leq)$
11 0 64 6
33 -1.0987 99 10
55 -3.1074 163 16
99 -2.6027 147 15
$49$ -5.1446 229 22
$99$ -12.5690 466 45
$999$ -71.6034 2357 229
$9999$ -1227.9235 39384 3816
$49999$ -152574.2718 4885724 473399
$99999$ -16325.0518 522817 50658
$999999$ -15030.9423 481377 46643
@@ Acknowledgements
The third-named author acknowledges the support from IISER Pune, Bhaskaracharya Pratishthana (Pune), and NISER Bhubaneswar and his work was also supported at NISER Bhubaneswar by the Department of Atomic Energy (DAE project number RIN-4001). The second-named author is grateful to COEP Technological University, Pune, for providing an excellent working environment. The second and third named authors would like to thank Kaneenika Sinha for her valuable input. We are thankful to the Harish-Chandra Research Institute, Prayagraj, for excellent hospitality, where the parts of the work were carried out. We are grateful to the anonymous referee for pointing out an error in Theorem 1 and consequent results.
Data availability
The authors declare that the manuscript has no associated data.